For each of the following equations, determine whether demand is elastic, inelastic, or unitary at given price
a. 4P = 100 - Q and P= $20
b. 20P = 1500 - Q and P=$5
c. P=50 - 0.1Q and P=$20
1
Expert's answer
2018-03-14T06:42:08-0400
For each of the following equations, determine whether demand is elastic, inelastic, or unitary at given price
a. 4P = 100 - Q, Q = 100 - 4P,
P = $20, so Q = 20,
Ed = -4*20/20 = -4, so demand is elastic.
b. 20P = 1500 - Q, Q = 1500 - 20P,
P = $5, so Q = 1400,
Ed = -20*5/1400 = -0.07, so demand is inelastic.
c. P = 50 - 0.1Q, Q = 500 - 10P,
P = $20, so Q = 300,
Ed = -10*20/300 = -0.67, so demand is inelastic.